# Algorithmic Decision-Making under Agents with Persistent Improvement

### v1

@@ 3 Improvement & Optimal Effort
Based on $C(\theta,r,x_{0})$, we can derive conditions under which agents have incentives to improve (Thm. 3.1).
@@
Thm. 3.1 identifies a condition under which agents have incentives to exert positive effort $k>0$. This condition depends on factors $\theta,r,x_{0}$ and can be fully characterized by the function $C:=C(\theta,r,x_{0})$.
@@
Although the analytical solution of the threshold $m$ is difficult to find, we can numerically solve $m\approx 0.3164$ as shown in App. H.1. In Fig. 2, we illustrate agent utilities $U$ as functions of effort under different $C$. The results show that only when $C<m$ (red curve), an agent can attain positive utility with effort $k>0$; when $C\geq m$ (green/yellow/blue curve), agents will not invest because the maximum utility is attained at $k=0$. Moreover, when $C<m$ (red curve), there is a unique optimal effort $k$ that maximizes the utility. These results are consistent with Thm. 3.1.
The condition in Thm. 3.1 further indicates the impacts of policy $\theta$, discounting factor $r$, and initial state $x_{0}$ on agent behavior. Specifically, agents only invest if $C(\theta,x_{0},r)<m$ holds. By fixing any two of $\theta,x_{0},r$, we can identify the domain of the third factor under which agents invest to improve. These results are summarized in Table 1 and verified in App. D. It shows that for any threshold $\theta$ and discounting factor $r$, agents only improve if their initial qualification profile is sufficiently similar to the ideal profile; the domain of $\theta$ also implies the best profile an agent with initial state $x_{0}$ can reach after exerting effort: if acceptance threshold $\theta$ is larger than the upper bound of $\theta$ given in Table 1, then agents will not have incentives to improve.
@@
The above results further suggest effective strategies that encourage agents to improve their qualifications, i.e., more agents are incentivized to improve if (i) the decision-maker’s acceptance threshold $\theta$ is lower; or (ii) the time it takes for agents to succeed after investments is shorter (smaller discounting factor $r$). Examples of both strategies in real applications are discussed in App. D, which further verify the effectiveness of our proposed model.
@@ 4 Decision-maker’s policy to incentivize improvement
Eq. (7) above demonstrates that the decision-maker aims to maximize the total improvement among the agent population, and its utility is a function of $\theta$. Since $f(x),x^{*}(\theta)$ are both continuous in $[0,1]$, utility $U_{d}(\theta)$ is also continuous. The following Thm. 4.1 further shows the existence of the optimal thresholds $\theta^{*}\in(0,1)$.
@@
To verify Thm. 4.1, we demonstrate the values of $U_{d}$ under situations where the agent population has different density functions $f$ and different discounting factors $r$. Specifically, we consider the uniform distribution and Beta distributions with different parameters. Fig. 3 shows $U_{d}(\theta)$ under different density functions $f$ and discounting factors $r$. The results illustrate that under these settings, $U_{d}$ is single-peaked and there is a unique $\theta^{*}\in(0,1)$ that is optimal and results in positive utility, which is consistent with Thm. 4.1. The figure also indicates the impact of $r$ on the optimal threshold: as $r$ increases, $\theta^{*}$ increases and the corresponding maximum utility decreases. As formally stated below in Corollary 4.2. We prove Thm. 4.1 and Corollary 4.2 in App. H.2.
@@
Importantly, the results of Thm. 4.1 show that the decision-maker can always find an optimal decision threshold $\theta^{*}$ (either numerically or using gradient methods depending on the density function $f$) to incentivize the largest improvement and promote social welfare in practice. While the above results all assume the decision-maker knows $r$ when determining $\theta$, we can relax this and provide a procedure to estimate $r$; this is included in App. G.
@@ 5 Impact of Manipulative Behavior
Model with both manipulation & improvement. We extend the model in Sec. 2 where agents after observing $\theta$ have an additional option to manipulate $x_{0}$ directly. If they choose to improve, they make a one-time effort $k\in[0,1]$ to acquire relevant knowledge and gradually improve their qualifications $q_{t}$ over time based on (1). If they choose to manipulate, they only increase $x_{t}$ at every round to fool the decision-maker without changing the actual profile $q_{t}$. Similar to the literature on strategic classification (Hardt et al., 2016a), the manipulation comes at the cost and the risk of being caught.
@@
Degree of manipulation. If agents choose to manipulate, they will increase $x_{t}$ at every round to fool the decision-maker, and they manipulate in a way that minimizes the manipulation cost and the risk of being detected. We make the following natural assumptions on $c$ and $P$:
@@
Agent’s best response. Suppose agents have full information about detection probability $P$ and discounting factor $r$, after observing the acceptance threshold $\theta$, they best respond by choosing the action (i.e., improvement/manipulation/do nothing) that maximizes their utilities, i.e., if $\widetilde{U}>\max_{k}{U}$, they choose to manipulate; otherwise, they improve by exerting optimal effort $k^{*}=\arg\max_{k}{U}$.
@@
Thm. 5.1 considers scenarios when the threshold is sufficiently high, and identifies conditions under which manipulation is preferred by agents in these settings. It shows agent behavior highly depends on the risk of manipulation (i.e., detection probability $P$). The specific values of $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ in Thm. 5.1 depend on $\theta$, $r$. In particular, $\widehat{P}$ increases as $r$ increases. Indeed, we can empirically find $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ and verify the theorem. These are illustrated in App. E and Sec. 7.
@@ 6 Forgetting Mechanism
Thm. 6.1 implies that convergence still holds when qualifications evolve with forgetting. Unlike the scenarios without forgetting where $q_{t}$ eventually converges to the ideal profile $d$ regardless of $k$ (Lemma 2.1), $q_{t}$ now converges to $d^{*}$, i.e., a profile between initial qualifications $q_{0}$ and ideal profile $d$, which is closer to $q_{0}$ with smaller investment $k$. It shows that if agents do not exert enough effort and the acquired knowledge is not sufficient, then they will not make satisfactory improvements.
Agent’s utility and improvement action. Denote agent utility under the forgetting mechanism as $\widehat{U}(k,\theta,r,x_{0})$. Unlike settings without forgetting where we can derive the exact time $H$ it takes for agents to be accepted and find utility $U$ ((5)), the analytical form of $\widehat{U}(k,\theta,r,x_{0})$ is not easy to derive. Nonetheless, we can still show that there exist scenarios under which agents have incentives to improve, even though the best attainable profile is a profile $d^{*}$ between initial $q_{0}$ and the ideal $d$.
@@
$\widehat{k}=\min\left(\frac{\widehat{x}_{0}^{2}}{2\widehat{x}_{0}^{2}+2\widehat{x}_{0}^{3}},\frac{x_{0}\cdot(x_{0}^{2}+x_{0}-\sqrt{x_{0}^{4}-x_{0}^{2}+1})}{2x_{0}^{2}+2x_{0}^{3}-1}\right)$
@@
Thm. 6.2 implies that there exists $(\theta,r)$ such that agents best respond by improving their qualifications, and the optimal effort is upper bounded by $\widehat{k}$. Indeed, we can numerically find the upper bound $\widehat{k}$ as a function $x_{0}$ (shown in Fig. 4). Because $\widehat{k}<0.35$ for all $x_{0}$, the improvement an agent can make under the forgetting mechanism is limited, suggesting that the agents may not improve to be qualified when the tasks are challenging enough.
Remark.
Under the forgetting mechanism, the actual effort invested by any agent is less than 0.35, and the qualifications $q_{t}$ converge to a profile $d^{*}$ that is between $q_{0}$ and $0.35\cdot d+0.65\cdot q_{0}$.
@@ 7 Experiments
Refer to captionRefer to caption
Refer to captionRefer to caption
We validate theoretical results by conducting experiments on Exam score (Kimmons, 2012) and FICO score (Reserve, 2007) dataset. For both datasets, scores serve as the agent’s initial similarity $x_{0}$, and we assume agents interact with a decision maker based on the Stackelberg game in Sec. 2. We first fit these scores with beta distributions, i.e., $x_{0}\sim\text{Beta}(v,w)$, and then use them to derive the followings:
@@
Exam Score Data. It is a synthetic dataset containing 1000 students’ exam scores on 3 subjects including math, reading, and writing (Kimmons, 2012). We first average over 3 subjects and normalize the averaged score to $[0,1]$. Then, we fit two beta distributions to the normalized scores of males and females and obtain $x_{0}\sim\text{Beta}(4.86,2.37),\text{Beta}(4.15,1.79)$ (see Fig. 9 in App. F).
@@
FICO Score Data. We adopt the data pre-processed by Hardt et al. (2016b), which contains CDF of credit scores of four racial groups (Caucasian, African American, Hispanic, Asian). For each group, we fit a Beta distribution and obtain four distributions: $\text{Beta}(1.11,0.97)$ for Caucasian, $\text{Beta}(0.91,3.84)$ for African American, $\text{Beta}(0.99,1.58)$ for Hispanic, $\text{Beta}(1.35,1.13)$ for Asian (see Fig. 10 in App. F). We only present the results for Caucasians and African Americans, while the results for Asian and Hispanic are shown in App. F.
@@ 8 Conclusions & Limitations
However, our theoretical results depend on the assumption that both agents and the decision-maker have perfect information about each other so that they always best respond. Extension to cases when each party only has partial or imperfect information is important. Moreover, these theorems are based on the qualification dynamics (1). Although a scenario when it does not hold is studied in Sec. 6, future works should also consider other variants tailored to specific applications to prevent negative outcomes.
Impact Statement
@@
Appendix A Pre-normalization
Though we focus on the similarity between $q_{t}$ and $d$, magnitudes of $q_{t}$ do matter in many practical situations. For example, if the decision-maker prefers students with balanced math and English skills and there are two “balanced" students, the decision-maker will certainly prefer the one with higher scores. Therefore, we propose a pre-normalization procedure to incorporate magnitude into account. Specifically, we add an additional dimension representing the unobservable “irrelevant attributes” to $q_{0}$ and obtain a $m+1$ dimensional complete qualification profile. Meanwhile, we add an additional dimension to the ideal qualification profile $d$ with 0 as its value; the new ideal profile becomes $[d;0]$. Then we can make the following natural assumption:
Assumption A.1.
Assumption A.1 has been supported by literature in machine learning Liu et al. (2022) and social science Holmstrom and Milgrom (1991). The “irrelevant" dimension demonstrates all other skills that belong to an agent but are not important to the decision. Therefore, competency in relevant/measurable attributes implies weakness in irrelevant/immeasurable attributes and the length of the complete qualification profile stays the same for all agents. With Assumption A.1 and the distribution of $q_{0}$ as $Q$, we formalize the pre-normalization procedure in Algorithm 1.
1:  $d=[d,0]$.
3:  for $i\in\{1,\ldots n\}$ do
5:     $q_{0}^{i}=\frac{[q_{0}^{i};z^{i}]^{T}}{K}\in[0,1]^{m+1}$.
Algorithm 1 Pre-normalization procedure
Appendix B Discussion and Generalization of (1)
More details on the dynamics in (1). In the main paper, we assume the influence of the initial effort $k$ is persistent and will enable $q_{t}$ changes gradually during each round. This is well-supported by the following examples:
@@ Appendix B Discussion and Generalization of (1)
Model generalization when agents can invest efforts at different time steps. We discuss how the model in the main paper can capture more complicated scenarios where agents repeatedly exert efforts multiple times until they reach the target. Each effort has persistent effects on improving the qualification as shown in Eqn. (11).
@@
Model generalization when $k$ diminishes with $t$. In the main paper, $k_{t}$ is always equal to $k$, demonstrating the effort has a consistent and persistent effect on the improvement of an individual. According to Lemma 2.1, the similarity $x_{t}$ approaches $1$ at an exponential rate. Thus, the case of $k_{t}>k$ is not interesting since the convergence is faster and it may not make sense in practice that the effort can be increasingly effective as time goes on. However, in reality, it may be possible that $k_{t}$ is decreasing. This is a “middle-point" case between the regular improvement in (1) and the forgetting mechanism (9), which may illustrate the “tiredness" when agents stick to improve. However, we can prove that when $k_{t}$ decreases linearly (i.e., $k_{t}=\Theta(\frac{k}{t})$), the similarity $x_{t}$ can only converge to $1$ at a speed $\Theta(t^{k})$.
Theorem B.1.
@@
We prove Thm. B.1 in App. H.5. Basically, this result illustrates that the agents will still improve to be qualified if $k_{t}$ decreases at a linear rate. Specifically, we can rewrite the (2) as:
@@
Appendix C Related Work
C.1 Strategic Manipulation
Though our work primarily lies in proposing a new model for improvement behaviors, the problem settings are also closely related to strategic classification problems Hardt et al. (2016a); Ben-Porat and Tennenholtz (2017); Dong et al. (2018); Braverman and Garg (2020); Sundaram et al. (2021); Jagadeesan et al. (2021); Ahmadi et al. (2021); Eilat et al. (2022); Horowitz and Rosenfeld (2023). Hardt et al. (2016a) formulated classification problems with strategic manipulation as a Stackelberg game with deterministic cost functions, where the decision maker optimizes classification accuracy based on individuals’ best responses. Afterwards, more sophisticated analytical frameworks were proposed Dong et al. (2018); Braverman and Garg (2020); Jagadeesan et al. (2021). Dong et al. (2018) proposed an online algorithm for strategic classification, and Braverman and Garg (2020) added randomness to strategic classifiers. On the other hand, Sundaram et al. (2021) analyzes the statistical learnability of strategic classification with an SVC classifier. Jagadeesan et al. (2021) relaxed the standard microfoundations assumption where individuals are perfectly rational to alternative microfoundations where a proportion of individuals may not be strategic, and proposed a noisy response model to tackle the new problem. Zhang et al. (2022) studied the setting where the decision maker and individuals only have knowledge of the feature distributions as random variables. Thus, the strategic manipulation corresponds to a distribution shift and its cost is also a random variable. Eilat et al. (2022) considered the setting where individual responses are dependent and the classifier is learned through graph neural networks.
C.2 Improvement
However, there are other literature considering improvement behaviorLiu et al. (2019); Rosenfeld et al. (2020); Shavit et al. (2020); Alon et al. (2020); Zhang et al. (2020); Chen et al. (2020); Kleinberg and Raghavan (2020); Bechavod et al. (2021); Ahmadi et al. (2022a, b); Raab and Liu (2021). Unlike strategic manipulation, improvement will incur a label change. Liu et al. (2019) studied the conditions where fairness interventions can promote improvement among individuals. Zhang et al. (2020) formulated the label change as a transition matrix where the transition probabilities are deterministic and difficult to estimate. Other works consider both behaviors at the same time. Kleinberg and Raghavan (2020) proposed a mechanism to incentivize individuals to invest in specific features where the individuals have a budget to invest strategically on all features including undesired ones. Their work inherited the classical settings of the Principal-agent model in economics but designed an incentivizing mechanism under a linear machine learning classifier. They modeled manipulation and improvement similarly (linear in efforts) and did not consider the persistent and delayed effects of improvement. By contrast, we first develop an fundamentally different dynamic model to characterize persistent and delayed improvements. Based on this model, we construct a Stackelberg game to model the interplay between agents and the decision-maker. Shavit et al. (2020) and Alon et al. (2020) introduced causal inference frameworks into strategic behaviors including manipulation and improvement. Chen et al. (2020) divided the features into immutable features, improvable features and manipulable features and explored linear classifiers which can prevent manipulation and encourage improvement. Jin et al. (2022) also focused on incentivizing improvement and proposed a subsidy mechanism to induce improvement actions and improve social well-being metrics. Barsotti et al. (2022) conducted several empirical experiments when both improvement and manipulation are possible where both actions incur a linear deterministic cost.
C.3 Recommendation Systems
Our work is also related to preference shifts and opinion dynamics in recommendation systems, which we refer to Castellano et al. (2009) as a comprehensive survey. Among the rich set of works, Dean and Morgenstern (2022); Gaitonde et al. (2021) proposed geometric models for opinion polarization and motivate our work.
Appendix D Illustration of Table 1

### v2

@@ 4 Improvement & Optimal Effort
Based on $C(\theta,r,x_{0})$, we can derive conditions under which agents have incentives to improve (Thm. 4.1).
@@
Thm. 4.1 identifies a condition under which agents have incentives to exert positive effort $k>0$. This condition depends on factors $\theta,r,x_{0}$ and can be fully characterized by the function $C:=C(\theta,r,x_{0})$.
@@
Although the analytical solution of the threshold $m$ is difficult to find, we can numerically solve $m\approx 0.3164$ as shown in App. F.1. In Fig. 2, we illustrate agent utilities $U$ as functions of effort under different $C$. The results show that only when $C<m$ (red curve), an agent can attain positive utility with effort $k>0$; when $C\geq m$ (green/yellow/blue curve), agents will not invest because the maximum utility is attained at $k=0$. Moreover, when $C<m$ (red curve), there is a unique optimal effort $k$ that maximizes the utility. These results are consistent with Thm. 4.1.
The condition in Thm. 4.1 further indicates the impacts of policy $\theta$, discounting factor $r$, and initial state $x_{0}$ on agent behavior. Specifically, agents only invest if $C(\theta,x_{0},r)<m$ holds. By fixing any two of $\theta,x_{0},r$, we can identify the domain of the third factor under which agents invest to improve. These results are summarized in Table 1 and verified in App. B. It shows that for any threshold $\theta$ and discounting factor $r$, agents only improve if their initial qualification profile is sufficiently similar to the ideal profile; the domain of $\theta$ also implies the best profile an agent with initial state $x_{0}$ can reach after exerting effort: if acceptance threshold $\theta$ is larger than the upper bound of $\theta$ given in Table 1, then agents will not have incentives to improve.
@@
The above results further suggest effective strategies that encourage agents to improve their qualifications, i.e., more agents are incentivized to improve if (i) the decision-maker’s acceptance threshold $\theta$ is lower; or (ii) the time it takes for agents to succeed after investments is shorter (smaller discounting factor $r$). Examples of both strategies in real applications are discussed in App. B, which further verify the effectiveness of our proposed model.
@@ 5 Decision-maker’s policy to incentivize improvement
Eq. (7) above demonstrates that the decision-maker aims to maximize the total improvement among the agent population, and its utility is a function of $\theta$. Since $f(x),x^{*}(\theta)$ are both continuous in $[0,1]$, utility $U_{d}(\theta)$ is also continuous. The following Thm. 5.1 further shows the existence of the optimal thresholds $\theta^{*}\in(0,1)$.
@@
To verify Thm. 5.1, we demonstrate the values of $U_{d}$ under situations where the agent population has different density functions $f$ and different discounting factors $r$. Specifically, we consider the uniform distribution and Beta distributions with different parameters. Fig. 3 shows $U_{d}(\theta)$ under different density functions $f$ and discounting factors $r$. The results illustrate that under these settings, $U_{d}$ is single-peaked and there is a unique $\theta^{*}\in(0,1)$ that is optimal and results in positive utility, which is consistent with Thm. 5.1. The figure also indicates the impact of $r$ on the optimal threshold: as $r$ increases, $\theta^{*}$ increases and the corresponding maximum utility decreases. As formally stated below in Corollary 5.2. We prove Thm. 5.1 and Corollary 5.2 in App. F.2.
@@
Importantly, the results of Thm. 5.1 show that the decision-maker can always find an optimal decision threshold $\theta^{*}$ (either numerically or using gradient methods depending on the density function $f$) to incentivize the largest improvement and promote social welfare in practice. While the above results all assume the decision-maker knows $r$ when determining $\theta$, we can relax this and provide a procedure to estimate $r$; this is included in App. E.
@@ 6 Impact of Manipulative Behavior
Model with both manipulation & improvement.
We extend the model in Sec. 3 where agents after observing $\theta$ have an additional option to manipulate $x_{0}$ directly. If they choose to improve, they make a one-time effort $k\in[0,1]$ to acquire relevant knowledge and gradually improve their qualifications $q_{t}$ over time based on (1). If they choose to manipulate, they only increase $x_{t}$ at every round to fool the decision-maker without changing the actual profile $q_{t}$. Similar to the literature on strategic classification (Hardt et al. 2016a), the manipulation comes at the cost and the risk of being caught.
@@
Degree of manipulation.
If agents choose to manipulate, they will increase $x_{t}$ at every round to fool the decision-maker, and they manipulate in a way that minimizes the manipulation cost and the risk of being detected. We make the following natural assumptions on $c$ and $P$:
@@
Agent’s best response.
Suppose agents have full information about detection probability $P$ and discounting factor $r$, after observing the acceptance threshold $\theta$, they best respond by choosing the action (i.e., improvement/manipulation/do nothing) that maximizes their utilities, i.e., if $\widetilde{U}>\max_{k}{U}$, they choose to manipulate; otherwise, they improve by exerting optimal effort $k^{*}=\arg\max_{k}{U}$.
@@
Thm. 6.1 considers scenarios when the threshold is sufficiently high, and identifies conditions under which manipulation is preferred by agents in these settings. It shows agent behavior highly depends on the risk of manipulation (i.e., detection probability $P$). The specific values of $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ in Thm. 6.1 depend on $\theta$, $r$. In particular, $\widehat{P}$ increases as $r$ increases. Indeed, we can empirically find $\widehat{P}$, $\widehat{x}$, $\widehat{x_{1}}$, $\widehat{x_{2}}$ and verify the theorem. These are illustrated in App. C and Sec. 8.
@@ 7 Forgetting Mechanism
Thm. 7.1 implies that convergence still holds when qualifications evolve with forgetting. Unlike the scenarios without forgetting where $q_{t}$ eventually converges to the ideal profile $d$ regardless of $k$ (Lemma 3.2), $q_{t}$ now converges to $d^{*}$, i.e., a profile between initial qualifications $q_{0}$ and ideal profile $d$, which is closer to $q_{0}$ with smaller investment $k$. It shows that if agents do not exert enough effort and the acquired knowledge is not sufficient, then they will not make satisfactory improvements.
Agent’s utility and improvement action.
Denote agent utility under the forgetting mechanism as $\widehat{U}(k,\theta,r,x_{0})$. Unlike settings without forgetting where we can derive the exact time $H$ it takes for agents to be accepted and find utility $U$ ((5)), the analytical form of $\widehat{U}(k,\theta,r,x_{0})$ is not easy to derive. Nonetheless, we can still show that there exist scenarios under which agents have incentives to improve, even though the best attainable profile is a profile $d^{*}$ between initial $q_{0}$ and the ideal $d$.
@@
$\widehat{k}=\min\left(\frac{\widehat{x}_{0}^{2}}{2\widehat{x}_{0}^{2}+2\widehat{x}_{0}^{3}},\frac{x_{0}\cdot(x_{0}^{2}+x_{0}-\sqrt{x_{0}^{4}-x_{0}^{2}+1})}{2x_{0}^{2}+2x_{0}^{3}-1}\right)$ (11)
@@
Thm. 7.2 implies that there exists $(\theta,r)$ such that agents best respond by improving their qualifications, and the optimal effort is upper bounded by $\widehat{k}$. Indeed, we can numerically find the upper bound $\widehat{k}$ as a function $x_{0}$ (shown in Fig. 4). Because $\widehat{k}<0.35$ for all $x_{0}$, the actual effort invested by any agent is less than 0.35, and the qualifications $q_{t}$ converge to a profile $d^{*}$ that is between $q_{0}$ and $0.35\cdot d+0.65\cdot q_{0}$. This means the improvement an agent can make under the forgetting mechanism may be limited, suggesting that the agents may not improve to be qualified when the tasks are challenging.
@@ 8 Experiments
We validate theoretical results by conducting experiments on Exam score (Kimmons 2012) and FICO score (Reserve 2007) dataset [^2] https://github.com/osu-srml/Alg-Persistent-Improvement. For both datasets, scores serve as the agent’s initial similarity $x_{0}$, and we assume agents interact with a decision maker based on the Stackelberg game in Sec. 3. We first fit these scores with beta distributions, i.e., $x_{0}\sim\text{Beta}(v,w)$, and then use them to derive the followings:
@@
Exam Score Data.
It is a synthetic dataset containing 1000 students’ exam scores on 3 subjects including math, reading, and writing (Kimmons 2012). We first average over 3 subjects and normalize the averaged score to $[0,1]$. Then, we fit two beta distributions to the normalized scores of males and females and obtain $x_{0}\sim\text{Beta}(4.86,2.37),\text{Beta}(4.15,1.79)$ (see Fig. 10 in App. D).
@@
FICO Score Data.
We adopt the data pre-processed by Hardt et al. (2016b), which contains CDF of credit scores of four racial groups (Caucasian, African American, Hispanic, Asian). For each group, we fit a Beta distribution and obtain four distributions: $\text{Beta}(1.11,0.97)$ for Caucasian, $\text{Beta}(0.91,3.84)$ for African American, $\text{Beta}(0.99,1.58)$ for Hispanic, $\text{Beta}(1.35,1.13)$ for Asian (see Fig. 11 in App. D). The results for Caucasians and African Americans are shown on the first column in Fig. 7, while the results for Asian and Hispanic are shown the second column.
@@ 9 Conclusions & Limitations
However, our theoretical results depend on the assumption that both agents and the decision-maker have perfect information about each other so that they always best respond. Extension to cases when each party only has partial or imperfect information is important. Moreover, these theorems are based on the qualification dynamics (1). Although a scenario when it does not hold is studied in Sec. 7, future works should also consider other variants tailored to specific applications to prevent negative outcomes. For example, they may consider the more complex situation where the decision-maker has a time-varying ideal profile $d_{t}$.
Ethical Consideration Statement
@@
Acknowledgement
This material is based upon work supported by the U.S. National Science Foundation under award IIS-2202699, by grants from the Ohio State University’s Translational Data Analytics Institute and College of Engineering Strategic Research Initiative.
Appendix A Discussion and Generalization of (1)
More details on the dynamics in (1).
In the main paper, we assume the influence of the initial effort $k$ is persistent and will enable $q_{t}$ changes gradually during each round. This is well-supported by the following examples:
@@ Appendix A Discussion and Generalization of (1)
Model generalization when agents can invest efforts at different time steps.
We discuss how the model in the main paper can capture more complicated scenarios where agents repeatedly exert efforts multiple times until they reach the target. Each effort has persistent effects on improving the qualification as shown in Eqn. (12).
@@
Model generalization when $k$ diminishes with $t$.
In the main paper, $k_{t}$ is always equal to $k$, demonstrating the effort has a consistent and persistent effect on the improvement of an individual. According to Lemma 3.2, the similarity $x_{t}$ approaches $1$ at an exponential rate. Thus, the case of $k_{t}>k$ is not interesting since the convergence is faster and it may not make sense in practice that the effort can be increasingly effective as time goes on. However, in reality, it may be possible that $k_{t}$ is decreasing. This is a “middle-point" case between the regular improvement in (1) and the forgetting mechanism (9), which may illustrate the “tiredness" when agents stick to improve. However, we can prove that when $k_{t}$ decreases linearly (i.e., $k_{t}=\Theta(\frac{k}{t})$), the similarity $x_{t}$ can only converge to $1$ at a speed $\Theta(t^{k})$.
Theorem A.1.
@@
We prove Thm. A.1 in App. F.6. Basically, this result illustrates that the agents will still improve to be qualified if $k_{t}$ decreases at a linear rate. Specifically, we can rewrite the (2) as:
@@
Appendix B Illustration of Table 1